A 2000 kg car traveling at a speed of 36 m/s skids to a halt on wet concrete where μk = 0.60. How long are the skid marks?

Show work!
Answer #1

m = 2000kg
u = 36m/s
mu = 0.6
deceleration = mu*g = 5.88m/s/s
s = 36^2/2*5.88 = 110.2 m